# Loss mechanisms of microwave frequency acoustic waves in thin film lithium niobate

### v1

@@ Abstract
Among many piezoelectric materials that are employed for SAW devices, lithium niobate (LN) stands out for its strong piezoelectricity and excellent mechanical and optical properties Slobodnik Jr et al. (1970); Lu and Gong (2021); Shao et al. (2022); Qiao et al. (2023). SAW with low propagation loss can be efficiently generated on LN, and high-performance SAW filters have been demonstrated on bulk LN. Recently, thin-film lithium niobate (TFLN) platforms, including lithium niobate on insulator (LNOI), lithium niobate on sapphire (LNOS), and suspended LN thin film, have emerged for many acoustic and optical applications. They support co-guided optical and acoustic modes with enhanced mode confinements and interaction strength. They enable efficient acousto-optic Li et al. (2023); Ye et al. (2025) and piezo-optomechanical devices Shen et al. (2020); Jiang et al. (2020). In addition, non-suspended TFLN substrates offer improved mechanical robustness and power-handling capability, making them particularly attractive for integrated SAW-photonic platforms.
@@
Figure 1: Device design and measurement principle. (a). An optical microscope image of an IDT pair fabricated on x-cut LNOI substrate. Two slightly chirped IDTs are fabricated along the LN $y$-axis with a delay length $L$. (b) Schematic illustration of acoustic propagation loss measurement scheme using a delay line. (c). Representative S-parameter measurement results of the device. Region I and II correspond to the excitation of Rayleigh and shear-horizontal acoustic modes, respectively. (d). Finite-element-method (FEM) simulations showing the displacement field distributions of the two acoustic modes.
@@
The acoustic propagation loss is measured using a delay-line method. A pair of interdigital transducers (IDTs) separated by a delay length $L$ is fabricated to excite and receive acoustic waves, as shown in Fig. 1(a). On samples A, B, and D, the acoustic waves propagate along the $y$-axis of the LN crystal, whereas on sample C, they propagate along the $x$-axis. A vector network analyzer (VNA) is used to apply drive signals of microwave frequencies to the first IDTs and measure the signals from the second IDTs to determine the device’s S-parameters. Frequency-domain measurement results can be transformed to the time domain for further analysis.
Figure 2: Acoustic loss measurement using a delay-line method. (a). Time-domain signals obtained by taking the inverse Fourier transform of the $S_{12}$ response within the 1.82-1.87 GHz frequency range. Each trace corresponds to an IDT pair with a different delay length $L$, varying from $400~\mu\mathrm{m}$ (bottom) to $800~\mu\mathrm{m}$ (top), with vertical offsets applied for clarity. The shaded wave packets indicate acoustic pulses arriving at the receiving IDT via direct transmission or multiple round-trip reflections. Cross and circle markers denote the peak amplitudes of the first and the second arrival wave packets, respectively. Measurements are performed at room temperature. (b). Acoustic group velocity extracted from a linear fit of the delay time of the first wave packet as a function of the corresponding delay length. (c). A 6.97 dB/mm propagation loss extracted from a linear fit of the measured amplitude ratio between the first and the second wave packets at different delay lengths $L$. (d). Frequency dependence of acoustic propagation loss, fitted with a quadratic function of frequency. The shaded regions in this figure and the following indicate the 95% confidence interval.
@@
, where $\eta_{1}$ ($\eta_{2}$), which will need independent calibration, are canceled out. We fabricate a set of devices with identical IDT designs and fabrication processes, but with varying delay lengths $L$. Assuming that the acoustic reflectivity $r_{1}$($r_{2}$) remain constant across these devices, the acoustic propagation loss $\alpha$ is extracted by fitting $\gamma(L)$ as a function of $L$.
@@
Figure 3: Temperature dependence of acoustic loss in TFLN platforms. (a) Acoustic propagation loss of different acoustic modes measured in LNOI and LNOS in the temperature range from 4 to 400 K. (b) The inverse quality factors extracted from measurements on the LNOI sample. (c) Zoom-in view of the acoustic propagation loss versus temperature in LNOS with a linear fit, where T is the temperature (K).
With the above measurement scheme, we study the frequency dependence of the acoustic loss. In dielectric crystals, a key intrinsic loss mechanism is due to interaction with thermal phonons in the medium Woodruff and Ehrenreich (1961). In the case $\omega\tau\ll 1$, where $\omega$ is the phonon angular frequency and $\tau$ is the thermal phonon relaxation rate, the acoustic loss is described by the Akhiezer damping model: $\alpha\propto T\omega^{2}$, where $\alpha$ is the propagation loss in dB/mm. When T decreases, or $\omega$ increases such that $\omega\tau\gg 1$, it is in the Landau and Rumer regime Pomerantz (2005), such that $\alpha\propto T\omega$. In bulk single-domain LN, the Akhiezer loss with frequency dependence $\alpha\propto\omega^{2}$ has been measured at room temperature over a broad frequency range extending to tens of gigahertz Bajak et al. (1981); Wen and Mayo (1966).
To obtain the frequency dependence of acoustic loss, we repeated measurements of the Rayleigh mode at different frequencies on LNOI. As shown in Fig. 2(d), the propagation loss reduces quadratically as frequency decreases from 2.96 GHz to 0.73 GHz, with a fitting function of $\alpha_{LNOI,f}=1.294\,f^{2}+3.71$ in the unit of dB/mm, where $f$ is the acoustic frequency in gigahertz. The frequency-dependent term of $1.294\,f^{2}$ is consistent with the result measured in crystalline LN, while the frequency-independent residue loss of 3.71 dB/mm indicates that an extra loss channel exists on the LNOI platform.
In addition to frequency dependence, Akhiezer damping is theoretically expected to vary linearly with temperature. In contrast, acoustic loss due to defect scattering can exhibit a strong non-monotonic temperature dependence Pohl et al. (2002). Defect scattering loss in amorphous materials is attributed to scattering by defect modes with a broad spectral distribution and can be well modeled with the tunneling model Phillips (1972). The ratio of the phonon wavelength $\lambda$ to the phonon mean free path $l$ of these studied amorphous materials Pohl et al. (2002) falls in the range of $10^{-3}-10^{-2}$.
Motivated by these distinct temperature-dependent loss behaviors, we investigated the temperature-dependent acoustic loss of all four platforms. Because the intrinsic mechanical properties of LN are only weakly dependent on temperature Smith and Welsh (1971), temperature-dependent measurements are effective to infer the dominant acoustic loss channels in each platform.
@@
Fig. 3(a) shows the temperature dependence of acoustic loss in LNOI and LNOS measured from 4 K to 400 K. In LNOS, the acoustic loss decreases with decreasing temperature; a zoom-in view of the data together with a linear fit is shown in Fig. 3(c), suggesting that the dominant loss mechanism could be Akhiezer damping. In contrast, the acoustic loss in LNOI exhibits a non-monotonic temperature dependence. For both Rayleigh and SH modes, the acoustic loss increases by 20–40 dB/mm as the temperature is lowered from room temperature to $\sim$90 K. As the temperature is further reduced, the acoustic loss decreases from its peak value but remains higher than the room temperature value.
@@
We further calculate the inverse quality factor $Q^{-1}$ for the LNOI sample using the measured propagation loss (in dB/mm) shown in Fig. 3(a). Here, the quality factor is defined as $Q=2\pi\,l/\lambda$, with $\lambda$ and $l$ denoting the acoustic wavelength and phonon mean free path, respectively. The inverse quality factor is related to acoustic loss as $\alpha=\frac{0.0546\,f}{Q\,v_{g}}$, where $f$ is the acoustic frequency and $v_{g}$ is the group velocity. The extracted $Q^{-1}$ is shown in Fig. 3(b) and exhibits a temperature dependence similar to that reported for amorphous SiO2 in a previous work Pohl et al. (2002). The magnitude of $Q^{-1}$ also falls within the characteristic range of glass Pohl et al. (2002), suggesting that the buried-oxide (BOX) layer in LNOI is a dominant loss channel for acoustic waves at low temperatures. Interestingly, as the temperature is increased from room temperature to 350-400 K, the acoustic loss reduces, supporting the interpretation that the frequency-independent residual loss originates from defect mode scattering in the BOX layer. At 400 K, the measured loss ($4.56\pm 0.33$ dB/mm) is consistent with the frequency-dependent term shown in Fig. 2(d) within experimental uncertainty, indicating that the loss contribution from the BOX layer becomes negligible at this temperature.
@@
We also analyze the temperature-dependent acoustic propagation loss in bulk LN and suspended LN thin film. As shown in Fig. 4, the propagation losses measured on these two platforms show a linear temperature dependence from 4 K to room temperature, similar to LNOS. Attributing to the Akhiezer damping, the acoustic propagation loss extrapolated to different frequencies is of the same order of magnitude as previously reported values Slobodnik Jr et al. (1970); Lu et al. (2020). The linear fit indicates that bulk and suspended LN exhibit similar temperature-dependent acoustic loss but differ slightly in the temperature-independent residual loss. The latter could be attributed to defect scattering in crystals or at the released interface.
@@
Compared with bulk LN, the acoustic propagation losses in TFLN platforms are higher by several dB/mm, even after excluding the loss associated with the BOX layer. To further investigate the origin of this excess loss, we examine the crystal quality and bonding interfaces of the TFLN platforms using scanning transmission electron microscopy (STEM). The STEM imaging was performed on an aberration-corrected Nion UltraSTEM electron microscope operated at 60 kV, using a semi-convergence angle of 32 mrad. High-angle annular dark-field (HAADF) images were acquired with semi-collection angles ranging from 80-200 mrad, using a probe current of 10 pA. Fig. 5 (a), (b) shows STEM images of the bonding interface of the LNOI substrate, showing a good structural quality with a well-resolved LN lattice at the interface. An approximately 6 nm contrast variation is observed at the interface, which requires further analysis to determine its chemical composition. A similar bonding interface condition of LNOS is shown in Fig. 5 (c). Amorphization is observed at the interface, resulting in a brighter contrast of approximately 2.7 nm in the LN layer and a 7 nm-thick dark layer in the sapphire substrate. The darker contrast is likely due to ion milling during focused ion beam–transmission electron microscopy (FIB–TEM) sample preparation, which induces oxygen vacancies.
@@
Although the transition layer at the bonding interface is only a few nanometers thick, crystal impurities in this region can introduce elastic disorder and defect-induced scattering, thereby providing a measurable loss channel for SAWs. This mechanism may account for the observed temperature-independent propagation loss of approximately 1.97 dB/mm for the fundamental asymmetric Lamb (A0) mode in suspended LN and that of 1.11 dB/mm for the SH mode in LNOS. It may also contribute to the Rayleigh-mode loss in LNOI as part of the observed quadratic frequency-dependent loss.
@@
In conclusion, we have systematically investigated acoustic attenuation mechanisms in thin-film lithium niobate (TFLN) platforms by comparing acoustic propagation loss across LNOI, LNOS, suspended LN thin films, and bulk LN over wide frequency and temperature ranges. We find that acoustic loss in LNOI exhibits a pronounced non-monotonic temperature dependence, which is characteristic of amorphous materials. The result suggests that the buried oxide layer is the dominant loss channel at low temperatures. At elevated temperatures, the loss associated with the buried oxide layer is suppressed, and the acoustic loss is governed primarily by Akhiezer damping in crystalline LN. In contrast, acoustic loss in LNOS, bulk LN, and suspended LN thin films is dominated by Akhiezer damping over the entire measured temperature range and decreases monotonically with decreasing temperature. These results indicate that acoustic loss in LNOI can be mitigated by operating at elevated temperatures or by replacing the buried oxide layer with a crystalline substrate. In addition, high-resolution HAADF-STEM imaging reveals nanoscale interfacial disorder at the bonding interface, which may introduce an additional loss channel in TFLN devices relative to bulk LN. These results clarify the physical origins of acoustic loss in TFLN platforms and provide important material and structural guidelines for minimizing acoustic loss in future phononic, acousto-optic, piezo-optomechanics, and quantum acoustic devices.

### v2

# Temperature-dependent acoustic loss at microwave frequencies in thin-film lithium niobate

@@ Abstract
Among many piezoelectric materials that are employed for SAW devices, lithium niobate (LN) stands out for its strong piezoelectricity and excellent mechanical and optical properties Slobodnik Jr et al. (1970); Lu and Gong (2021); Shao et al. (2022); Qiao et al. (2023). SAWs with low propagation loss can be efficiently generated on LN, and high-performance SAW filters have been demonstrated on bulk LN. Recently, thin-film lithium niobate (TFLN) platforms, including lithium niobate on insulator (LNOI), lithium niobate on sapphire (LNOS), and suspended LN thin film, have emerged for many acoustic and optical applications. They support co-guided optical and acoustic modes with enhanced mode confinements and interaction strength. They enable efficient acousto-optic Li et al. (2023); Ye et al. (2025) and piezo-optomechanical devices Shen et al. (2020); Jiang et al. (2020). In addition, non-suspended TFLN substrates offer improved mechanical robustness and power-handling capability, making them particularly attractive for integrated SAW-photonic platforms.
@@
Figure 1: Device design and measurement principles. (a). An optical microscope image of an IDT pair fabricated on X-cut LNOI substrate. Two slightly chirped IDTs are fabricated along the LN $y$-axis with a delay length $L$. The blue triangles indicate the probing locations using a pair of GSG probes. (b) Schematic illustration of the acoustic propagation loss measurement scheme using a delay line. (c). Representative S-parameter measurement results of the device. Regions I and II correspond to the excitation of Rayleigh and shear-horizontal acoustic modes, respectively. (d). Finite-element-method (FEM) simulations showing the displacement field distributions of the two acoustic modes. (e) Schematic illustrations of the material stacks of the samples studied in this work: (A) LNOI, (B) LNOS, (C) bulk LN, and (D) free-standing LN thin film.
@@
The acoustic propagation loss is measured using a delay-line method. A pair of interdigital transducers (IDTs) separated by a delay length $L$ is fabricated to excite and receive acoustic waves, as shown in Fig. 1(a) (see Supplementary Information S1 for the fabrication process). On samples A, B, and D, the acoustic waves propagate along the Y-axis of the LN crystal, whereas on sample C, they propagate along the X-axis. A vector network analyzer (VNA) is used to excite one IDT and measure the response at the other IDT to obtain the device S-parameters (see Supplementary Information S2 for the calibration and measurement process). Frequency-domain measurement results can be transformed to the time domain for further analysis.
Figure 2: Acoustic loss measurement using a delay-line method. (a). Time-domain signals obtained by taking the inverse Fourier transform of the $S_{21}$ response within the 1.82-1.87 GHz frequency range. Each trace corresponds to an IDT pair with a different delay length $L$, varying from $400~\mu\mathrm{m}$ (bottom) to $800~\mu\mathrm{m}$ (top), with vertical offsets applied for clarity. The shaded wave packets indicate acoustic pulses arriving at the receiving IDT via direct transmission or multiple round-trip reflections. Cross and circle markers denote the peak amplitudes of the first and the second arrival wave packets, respectively. Measurements are performed at room temperature. (b). Acoustic group velocity extracted from a linear fit of the delay time of the first wave packet as a function of the corresponding delay length. (c). A 6.97 dB/mm propagation loss extracted from a linear fit of the measured amplitude ratio between the first and the second wave packets at different delay lengths $L$. (d). Frequency dependence of acoustic propagation loss, fitted with a quadratic function of frequency. The shaded regions in this figure and the following indicate the 95% confidence interval, unless otherwise indicated.
@@
, where $\eta_{1}$ ($\eta_{2}$), which will need independent calibration, are canceled out. We fabricate a set of devices with identical IDT designs and fabrication processes, but with varying delay lengths $L$. Assuming that the acoustic reflectivities $r_{1}$ and $r_{2}$ remain the same across these devices, the acoustic propagation loss $\alpha$ is extracted by fitting $\gamma(L)$ as a function of $L$. It should be noted that the acoustic loss is extracted from free-propagating delay-line measurements without lateral confinement. As a result, the reported loss includes not only intrinsic propagation loss but also additional contributions arising from the measurement configuration, associated with beam walk-off and divergence, which reduce the overlap between the acoustic field and the receiving transducer. A detailed evaluation of these effects is provided in Supplementary Information S3.
@@
Figure 3: Temperature dependence of acoustic loss in TFLN platforms. (a) Acoustic propagation loss of different acoustic modes measured in LNOI and LNOS in the temperature range from 4 to 400 K. (b) The temperature-dependent quality factors extracted from measured acoustic loss in (a), acoustic frequencies, and group velocities of each mode.
With the above measurement scheme, we study the frequency dependence of the acoustic loss. In dielectric crystals, a key intrinsic loss mechanism is due to interaction with thermal phonons in the medium Woodruff and Ehrenreich (1961). In the case $\omega\tau\ll 1$, where $\omega$ is the phonon angular frequency and $\tau$ is the thermal phonon relaxation rate, the acoustic loss is described by the Akhiezer damping model: $\alpha\propto T\omega^{2}$, where $\alpha$ is the propagation loss in dB/mm. When T decreases, or $\omega$ increases such that $\omega\tau\gg 1$, it is in the Landau and Rumer regime Pomerantz (2005), such that $\alpha\propto T\omega$. In bulk single-domain LN, acoustic loss at room temperature has been reported to follow the Akhiezer loss mechanism with frequency dependence $\alpha\propto\omega^{2}$ up to frequencies approaching tens of gigahertz Bajak et al. (1981); Wen and Mayo (1966).
To obtain the frequency dependence of acoustic loss, we repeated measurements of the Rayleigh mode at different frequencies on LNOI. As shown in Fig. 2(d), the propagation loss reduces quadratically as frequency decreases from 2.96 GHz to 0.73 GHz, with a fitting function of $\alpha_{LNOI,f}=1.294\,f^{2}+3.71$ in the unit of dB/mm, where $f$ is the acoustic frequency in gigahertz. The frequency-dependent term of $1.294\,f^{2}$ is consistent with the result measured in crystalline LN, while the frequency-independent residue loss of 3.71 dB/mm indicates that extra loss channels may exist on the LNOI platform.
In addition to frequency dependence, Akhiezer damping is theoretically expected to vary linearly with temperature. In contrast, acoustic loss due to defect scattering can exhibit a strong non-monotonic temperature dependence Pohl et al. (2002). Defect scattering loss in amorphous materials is attributed to scattering by defect modes with a broad spectral distribution and can be well modeled with the tunneling model Phillips (1972). The ratio of the phonon wavelength $\lambda$ to the phonon mean free path $l$ of these studied amorphous materials Pohl et al. (2002) falls in the range of $10^{-3}-10^{-2}$ at low temperature.
Motivated by these distinct temperature-dependent loss behaviors, we investigated the acoustic loss of all four platforms as a function of temperature. Because the intrinsic mechanical properties of LN are only weakly dependent on temperature Smith and Welsh (1971), temperature-dependent measurements are effective to infer the dominant acoustic loss channels in each platform.
@@
Fig. 3(a) shows the temperature dependence of acoustic loss in LNOI and LNOS measured from 4 K to 400 K. In LNOS, the acoustic loss decreases monotonically with decreasing temperature. The acoustic loss in LNOS exhibits an approximately linear dependence on temperature, consistent with Akhiezer damping. In contrast, the acoustic loss in LNOI exhibits a non-monotonic temperature dependence. For both Rayleigh and SH modes, the acoustic loss increases by 20–40 dB/mm as the temperature is lowered from room temperature to $\sim$90 K. As the temperature is further reduced, the acoustic loss decreases from its peak value but remains higher than the room temperature value.
@@
We further calculate the quality factor $Q$ for the LNOI sample using the measured propagation loss (in dB/mm) shown in Fig. 3(a). Here, the quality factor is defined as $Q=2\pi\,l/\lambda$, with $\lambda$ and $l$ denoting the acoustic wavelength and phonon mean free path, respectively. The quality factor is related to acoustic loss as $Q=\frac{0.0546\,f}{\alpha\,v_{g}}$, where $f$ is the acoustic frequency and $v_{g}$ is the group velocity. The extracted $Q$ factors are shown in Fig. 3(b) and exhibit a temperature dependence similar to that reported for amorphous SiO2 in previous work Pohl et al. (2002). Interestingly, as the temperature is increased from room temperature to 350-400 K, the acoustic loss decreases. At 400 K, the measured loss ($4.56\pm 0.33$ dB/mm) agrees with the frequency-dependent term shown in Fig. 2(d) within experimental uncertainty, suggesting that the frequency-independent loss channel may be suppressed at this temperature.
@@
We also analyze the temperature-dependent acoustic propagation loss in bulk LN and suspended LN thin film. As shown in Fig. 4, the propagation losses measured on these two platforms show a linear temperature dependence from 4 K to room temperature, similar to LNOS. This behavior is consistent with Akhiezer damping. The acoustic propagation loss extrapolated to different frequencies is of the same order of magnitude as previously reported values Slobodnik Jr et al. (1970); Lu et al. (2020). The linear fit indicates that bulk and suspended LN exhibit similar temperature-dependent acoustic loss but differ slightly in the temperature-independent residual loss. The latter may arise from defect scattering in the crystal, scattering at the released interface, or strain variations induced during the release process.
@@
Compared with bulk LN, the acoustic propagation losses in TFLN platforms are higher by several dB/mm. To further investigate the origin of this excess loss, we examine the crystal quality and bonding interfaces of the TFLN platforms using scanning transmission electron microscopy (STEM). The STEM imaging was performed on an aberration-corrected Nion UltraSTEM electron microscope operated at 60 kV, using a semi-convergence angle of 32 mrad. High-angle annular dark-field (HAADF) images were acquired with semi-collection angles ranging from 80-200 mrad, using a probe current of 10 pA. Fig. 5(a), (b) shows STEM images of the bonding interface of the LNOI substrate, showing a good structural quality with a well-resolved LN lattice at the interface. An approximately 6 nm contrast variation is observed at the interface, which requires further analysis to determine its chemical composition. A similar bonding interface condition of LNOS is shown in Fig. 5(c). Amorphization is observed at the interface, resulting in a brighter contrast of approximately 2.7 nm in the LN layer and a 7 nm-thick dark layer in the sapphire substrate. The darker contrast is likely due to ion milling during focused ion beam–transmission electron microscopy (FIB–TEM) sample preparation, which induces oxygen vacancies.
@@
Although the transition layer at the bonding interface is only a few nanometers thick, crystal impurities in this region can introduce elastic disorder and defect-induced scattering, thereby providing a measurable loss channel for SAWs. This mechanism may account for the observed temperature-independent propagation loss of approximately 1.97 dB/mm for the fundamental asymmetric Lamb (A0) mode in suspended LN and that of 1.11 dB/mm for the SH mode in LNOS. It may also contribute to the acoustic loss in LNOI.
In conclusion, we have systematically investigated acoustic loss mechanisms in thin-film lithium niobate (TFLN) platforms by comparing acoustic propagation loss across LNOI, LNOS, suspended LN thin films, and bulk LN over wide frequency and temperature ranges. We find that acoustic loss in LNOI exhibits a pronounced non-monotonic temperature dependence, which may arise from the amorphous BOX layer or the distinct multilayer structure compared to the other three platforms. At elevated temperatures, the acoustic loss in LNOI is suppressed. In contrast, acoustic loss in LNOS, bulk LN, and suspended LN thin films exhibits a monotonic decrease with decreasing temperature over the entire measured range, consistent with Akhiezer damping. The acoustic loss extracted from the free-propagating delay line represents an effective loss and may deviate from the intrinsic material loss due to acoustic walk-off and beam divergence. These effects can be mitigated using waveguide-based delay lines. However, this contribution does not affect the observed trend for each platform or the main conclusions of this work.
@@
In addition, high-resolution HAADF-STEM imaging reveals nanoscale interfacial disorder at the bonding interface, which may introduce an additional loss channel in TFLN devices relative to bulk LN. These results clarify the physical origins of acoustic loss in TFLN platforms and provide important material and structural guidelines for minimizing acoustic loss in future phononic, acousto-optic, piezo-optomechanics, and quantum acoustic devices.